OpenStudy (anonymous):

Given the function f(x) = 2(x + 8), find x if f(x) = 12 2 40 −2 14

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

simply put x=12 into the equation 2(12+8) 2 x 20 = 40

OpenStudy (accessdenied):

It is if f(x) = 12, not x = 12. So, you would substitute in f(x) = 12 and solve for x.

OpenStudy (accessdenied):

40 would not be correct.

OpenStudy (anonymous):

2(x+8) = 12 => x+8 = 6 => x = -2